# PIP-Loco: A Proprioceptive Infinite Horizon Planning Framework for Quadrupedal Robot Locomotion

### v2

@@ I Introduction
Method Observation History Size Adaptation Module Internal Model Multi-terrain Planning
@@
On the other hand, model-free reinforcement learning (RL) methods have produced more robust and adaptive controllers compared to MPC methods. Such RL-based controllers have demonstrated significant potential for enabling quadruped locomotion across diverse terrains, including flat surfaces, stairs, slopes, and slippery environments [10, 11, 12]. The introduction of massively parallel RL agent training in IsaacGym [13] has further accelerated the development of RL controllers for quadrupeds. By leveraging this parallelism, [6] successfully designed a controller capable of adjusting various gait parameters, such as body height, step frequency, and step height, using Raibert’s heuristic. Additionally, the works of [14, 15, 16] have showcased RL’s ability to enable quadrupeds to perform parkour-like maneuvers. However, despite these advancements, several challenges still need to be solved in designing RL controllers, particularly regarding the robustness to sensor noise, interpretability of decision-making, and a mechanism to satisfy constraints online.
As mentioned earlier, RL-based quadruped controllers still are not robust enough to noise from hardware observations and suffer from a sim-to-real transfer gap. [9] proposed HIM-Loco that addresses noisy hardware observations by employing an internal model, drawing inspiration from [8]’s use of the asymmetric actor-critic framework. HIM-Loco demonstrated robustness across various terrains, such as sandy slopes, stairs, and bushes. However, their approach only considered one-step latent state prediction within the internal model and did not explore extending this model further. Training across diverse terrains simultaneously also often leads to policies that give generalized yet suboptimal actions. During deployment on a specific terrain, these actions can be improved through planning to achieve a more appropriate local optimum. Thus, taking inspiration from MPC and Dreamer World Models [17], in this paper, we aim to advance the internal model techniques and address challenges like constraint satisfaction and interpretability of behaviors by integrating planning mechanisms.
@@
We propose a proprioceptive infinite-horizon planning framework for quadruped locomotion, termed PIP-Loco (see Fig. 1). While training, we learn an expert policy (actor) and an internal model (comprising a velocity estimator and a Dreamer module) in a co-dependent manner. The Dreamer module includes a policy and value function cloned from the expert actor and privileged value function (critic), respectively, a reward function and a dynamics model to provide future states. In deployment, the Dreamer module solves an infinite-horizon optimization problem in an MPC fashion. This planning mechanism facilitates the filtering and adaptation of actions and velocity commands that might otherwise violate constraints, leading to robust quadruped locomotion. Table I compares our work with the state-of-the-art. The following are the broad contributions of our work:
@@ II-B4 Internal Model
Hence, in this work, we introduce an internal model that incorporates a velocity estimator, denoted as $v_{\phi}:\mathcal{O}^{M}\mapsto\mathbb{R}^{3}$, alongside three distinct variations of the Dreamer module. Each of these variants includes a dynamics model ($d_{\theta}$), a cloned policy model ($\pi_{\theta}$), a reward model ($r_{\theta}$), a value function ($V_{\theta}$), and an optional encoder ($\alpha_{\theta}$). While these variants differ in their architectural design, they share the same core components. Our motivation for studying these variants stems from the need to identify an effective "dreaming" process within the Dreamer module that can be successfully applied to solve an MPC problem during deployment. Below, we provide a detailed description of each variant, followed by an evaluation of their performance in Section III-A2.
@@ II-B5 Asymmetric Actor Critic
Many previous works ([7, 11]) have used adaptation modules to explicitly learn to encode privileged observations through regression. However, these adaptation modules typically require two separate training phases and are limited by approximation errors. In contrast, with our framework, the asymmetric actor-critic implicitly encourages the internal model outputs, which are fed to the expert actor, to encode information from privileged observations and enable dreaming.
The expert actor ($\pi_{e}$) receives the following inputs: the current observation ($o_{t}$), the current velocity estimate ($\hat{v}_{t}$), and the output of the Dreamer module ($y_{t}$). On the other hand, the privileged critic ($V_{e}$) takes as input both the current observation ($o_{t}$) and the privileged observation ($o^{e}_{t}$), which is obtained from the simulator. Privileged information includes the height field of the surface around the robot, external forces acting on the robot, the robot’s linear velocity, and other similar data.
A key distinction of our training approach, as shown in Fig. 2(a), is that the Dreamer module and the expert actor are interdependent, as both are learned simultaneously in a single phase. This co-dependence exists because the quality of $y_{t}$ relies on the actions predicted by $\pi_{\theta}$, which, in turn, depends on the performance of $\pi_{e}$.
@@ II-C Deployment
As discussed earlier, an online planner is essential for filtering out unsafe or suboptimal actions generated by the RL policy. It also enhances the interpretability of the robot’s behavior by providing insights into the agent’s dreaming process. To achieve this, we incorporate infinite-horizon planning inspired by techniques such as [21, 22, 23]. Our approach as shown in Fig. 2(b), involves solving the following optimization problem to obtain the optimal action sequence $a^{*}_{t:t+H}$ and a safe twist command $\nu^{*}_{cmd}:=(\text{v}^{cmd}_{xy},\omega^{cmd}_{z})$ at the current observation $o_{t}$,
@@
In this formulation, $\hat{o}_{t}$ denotes the observation where components corresponding to the velocity command in $o_{t}$ are replaced with the optimization variable $\nu_{cmd}$. The parameter $\gamma$ represents the discount factor, and $c_{k}(\cdot)\leq b_{k}$ specifies the constraints that must be satisfied at each step $k$.
@@
Solving the optimization problem in (3) is NP-hard. Various online optimization techniques, including sampling-based approaches that offer speed advantages over gradient or Hessian-based methods, can be used. In this work, we adopt a simple constrained Model Predictive Path Integral (MPPI) [24] approach, as detailed in Algorithm 1, with its hyperparameters listed in Table II. However, our framework is generic and can accommodate various other Model Predictive Control (MPC) algorithms such as those shown in [25, 26, 27], with each having its pros and cons.
We choose to use the Dreamer of NLM variant in Algorithm 1 due to its lower complexity and reduced inference cost compared to PLM and FLM. Recall that, unlike other models, NLM does not rely on any history of observations or latent spaces, which is the reason for its simplicity in implementation. Additionally, planning in latent spaces can compromise interpretability and limit our ability to enforce domain-specific constraints. In the next section, we comprehensively discuss our results with PIP-Loco.
@@ III Results
This section details the simulation and hardware experiments with the PIP-Loco. We also present ablation studies on the internal model to validate the importance of a $H$ step Dreamer. For training, we utilized an open-source environment for the Unitree Go1 quadruped robot, which is based on Nvidia’s Isaac Gym simulator [28] [13], and implemented all neural networks using the PyTorch framework [29]. The training process was executed on a desktop system with an AMD Ryzen Threadripper Pro 5995WX processor (128 cores) running at 2.7 GHz, 64 GB of RAM, and an NVIDIA RTX 3080 GPU. For real-time planning, we deployed an onboard computer using Jax [30], achieving an inference frequency of 300-500 Hz.
TABLE III: Performance comparison across different methods and noise levels after 2k iterations for 5 seeds
@@ III-A Abalations and Experiments
We conducted parallel training of 4096 robots across various terrains, employing domain randomization and a curriculum learning strategy. The PPO algorithm [19] used a clipping range of 0.2, a generalized advantage estimation factor of 0.95, and a discount factor of 0.99. All networks were optimized using the Adam optimizer [31] with a learning rate of 0.001.
Figure 4: Simulation results for a 10-second run with extreme command velocities: Top-left compares roll and pitch with and without planning. Top-right shows optimized velocity commands from Algorithm 1. The bottom panel shows joint angle variations, with planning having lower peaks near constraint bounds.
@@ III-A2 What Constitutes a Good Internal Model?
As discussed in [9], an internal model was introduced to simulate a single-step disturbance response in latent space. This work extends this approach by incorporating dreaming of future observations over $H$ steps. Figure 3 illustrates the reward curves for both the model proposed in [9] and our proposed internal models. Additionally, we investigate the influence of the prediction horizon $H$, the model’s robustness to noise, and the role of latent states within our framework. The corresponding results are summarized in Tables III(a) and III(b). Our experimental results suggest that the inclusion of latent states does not improve the robustness of the expert actor $\pi_{e}$ in the presence of training noise. However, leveraging future observation trajectories enables better exploration of gaits.
@@ III-A3 Advantages of Planning
We conducted an experiment to assess the robot’s ability to track extreme velocity commands, which can result in unstable behaviors. Fig. 4 presents a comparative analysis between the cloned policy $\pi_{\theta}$, which operates without planning, and Algorithm 1, which incorporates planning. The experiment aimed to track a target velocity of $\text{v}_{x}^{cmd}=2.0$\ \mathrm{m}\mathrm{/}\mathrm{s}$,\text{v}_{y}^{cmd}=2.0$\ \mathrm{m}\mathrm{/}\mathrm{s}$,\omega_{z}^{cmd}=0.5$\ \mathrm{rad}\mathrm{/}\mathrm{s}$$. The planning mechanism in PIP-Loco successfully optimized actions and velocity commands to minimize orientation deviations and maintain base stability while ensuring that joint positions have lower peaks that hit the constraint limits. As a future work, we aim to integrate obstacle avoidance and navigation-related constraints in the planner. The experimental results clearly demonstrate the need for planning in online adaptability. It is important to note that by integrating planning, we could add/remove constraints and reformulate reward/cost functions in real time without the need for retraining. This flexibility is a key characteristic of MPC approaches, which has been notably absent in RL methods for legged locomotion.
@@ IV Conclusion
We introduce PIP-Loco, a novel proprioceptive infinite-horizon planning framework for quadrupedal robot locomotion. Our approach overcomes the limitations of RL-based controllers in terms of interpretability and safety while addressing the task complexity and finite-horizon issue typical of classical MPC methods. By integrating our proposed internal model, PIP-Loco enhances the exploration of effective locomotion behaviors during training and solves an MPC problem at deployment, balancing safety with interpretability. The framework’s effectiveness is validated through ablation studies, which highlight the internal model’s quality and the importance of planning. Additionally, extensive testing in multi-terrain scenarios, both in simulation and on hardware, confirms its robustness and real-world potential. Future work will incorporate advancements in perceptive locomotion [15] and address compounding errors in dynamics predictions.

### v3

@@ I Introduction
Method Observation History Size Adaptation Module Internal Model Multi-terrain Planning (Deployment)
@@
On the other hand, model-free reinforcement learning (RL) methods have produced more robust and adaptive controllers compared to MPC methods. Such RL-based controllers have demonstrated significant potential for enabling quadruped locomotion across diverse terrains, including flat surfaces, stairs, slopes, and slippery environments [10, 11, 12]. The introduction of massively parallel RL agent training in IsaacGym [13] has further accelerated the development of RL controllers for quadrupeds. By leveraging this parallelism, [6] successfully designed a controller capable of adjusting various gait parameters, such as body height, step frequency, and step height, using Raibert’s heuristic. Additionally, the works of [14, 15, 16] have showcased RL’s ability to enable quadrupeds to perform parkour-like maneuvers.
Despite these advancements, several challenges still need to be solved when designing RL controllers. A critical issue is the reliable observability of essential states, such as the robot’s linear velocity, which affects both model-free and model-based approaches that typically combine state estimators to mitigate sensor noise. For this, [9] proposed HIM-Loco that addresses noisy hardware observations by employing an internal model, drawing inspiration from [8] ’s use of the asymmetric actor-critic framework. HIM-Loco demonstrated robustness across various terrains, such as sandy slopes, stairs, and bushes. However, their approach only considered one-step latent state prediction within the internal model and did not explore extending this model further. Training across diverse terrains simultaneously also often leads to policies that give generalized yet suboptimal actions. During deployment on a specific terrain, these actions can be improved through planning to achieve a more appropriate local optimum. Thus, taking inspiration from MPC and Dreamer World Models [17, 18], in this paper, we aim to advance the internal model techniques and address challenges like constraint satisfaction and interpretability of behaviors by integrating planning mechanisms.
@@
We propose a proprioceptive infinite-horizon planning framework for quadruped locomotion, termed PIP-Loco (see Fig. 1). While training, we learn an expert policy (actor) and an internal model (comprising a velocity estimator and a Dreamer module) in a co-dependent manner. The Dreamer module includes a policy and value function cloned from the expert actor and privileged value function (critic), respectively, a reward function and a dynamics model to provide future states. In deployment, the Dreamer module solves an infinite-horizon optimization problem in an MPC fashion. This planning mechanism facilitates the filtering and adaptation of actions and velocity commands that might otherwise violate constraints, leading to robust quadruped locomotion. Table I compares our work with the state-of-the-art. While we analyze multiple variants of the Dreamer module to assess the impact of increasingly complex internal models, our deployment relies on the simplest No-Latent Module (NLM) variant (see II-B4). The following are the broad contributions of our work:
@@
The benefits of our proposed framework are threefold. The first is overcoming limited observability in real-world settings through an internal model. The second is providing interpretability by leveraging the Dreamer module to simulate future trajectories, allowing for inspection and correction of predicted robot behaviors. Lastly, enabling the incorporation of hard constraints, such as joint limits or body stability conditions, directly into the planning process during deployment, ensuring safety and feasibility.
@@ II-B4 Internal Model
Hence, in this work, we introduce an internal model that incorporates a velocity estimator, denoted as $v_{\phi}:\mathcal{O}^{M}\mapsto\mathbb{R}^{3}$, alongside three distinct variations of the Dreamer module. Each of these variants includes a dynamics model ($d_{\theta}$), a cloned policy model ($\pi_{\theta}$), a reward model ($r_{\theta}$), a value function ($V_{\theta}$), and an optional encoder ($\alpha_{\theta}$). While these variants differ in architectural design, they share the same core components. Our motivation for studying these variants stems from the need to identify an effective "dreaming" process within the Dreamer module that can be successfully applied to solve an MPC problem during deployment. Below, we provide a detailed description of each variant, followed by an evaluation of their performance in Section III-A2.
@@ II-B5 Asymmetric Actor Critic
Many previous works ([7, 11]) have used adaptation modules to explicitly learn to encode privileged observations through regression. However, these adaptation modules typically require two separate training phases and are limited by approximation errors. In contrast, with our framework, the asymmetric actor-critic implicitly encourages the internal model outputs fed to the expert actor to encode information from privileged observations and enable dreaming.
The expert actor ($\pi_{e}$) receives the following inputs: the current observation ($o_{t}$), the current velocity estimate ($\hat{v}_{t}$), and the output of the Dreamer module ($y_{t}$). In our work, which builds on the HIM-Loco framework [9], the expert actor is trained to integrate these inputs effectively. On the other hand, the privileged critic ($V_{e}$) takes as input both the current observation ($o_{t}$) and the privileged observation ($o^{e}_{t}$) obtained from the simulator. This privileged observation includes information such as the height field of the surface around the robot, external forces acting on the robot, and the robot’s linear velocity, among other data.
As shown in Fig. 2(a), a key distinction of our training approach is that the Dreamer module and the expert actor are interdependent, as both are learned simultaneously in a single phase. This co-dependence exists because the quality of $y_{t}$ relies on the actions predicted by $\pi_{\theta}$, which, in turn, depends on the performance of $\pi_{e}$.
@@ II-C Deployment
As discussed earlier, an online planner is essential for filtering out unsafe or suboptimal actions generated by the RL policy. It also enhances the interpretability of the robot’s behavior by providing insights into the agent’s dreaming process. To achieve this, we incorporate infinite-horizon planning inspired by techniques such as [22, 23, 24]. Our approach, as shown in Fig. 2(b), involves solving the following optimization problem to obtain the optimal action sequence $a^{*}_{t:t+H}$ and a safe twist command $\nu^{*}_{cmd}:=(\text{v}^{cmd}_{xy},\omega^{cmd}_{z})$ at the current observation $o_{t}$,
@@
In this formulation, $\hat{o}_{t}$ denotes the observation where components corresponding to the velocity command in $o_{t}$ are replaced with the optimization variable $\nu_{cmd}$. $\nu_{cmd}$ may vary for each instance of $k$; however, for the sake of simplicity, it is assumed to remain consistent throughout the planning horizon. The parameter $\gamma$ represents the discount factor, and $c_{k}(\cdot)\leq b_{k}$ specifies the constraints that must be satisfied at each step $k$.
@@
Solving the optimization problem in (3) is NP-hard. To tackle this, there are various online optimization techniques, like sampling-based approaches, which offer speed advantages over gradient or Hessian-based methods. In this work, we adopt a simple constrained Model Predictive Path Integral (MPPI) [25] approach, as detailed in Algorithm 1, with its hyperparameters listed in Table II. However, our framework is generic and can accommodate various other Model Predictive Control (MPC) algorithms such as those shown in [26, 27, 28], with each having its pros and cons.
As stated earlier, we choose to use the Dreamer of NLM variant in Algorithm 1 due to its lower complexity and reduced inference cost compared to PLM and FLM. Recall that, unlike other models, NLM does not rely on any history of observations or latent spaces, which is the reason for its simplicity in implementation. Additionally, planning in latent spaces can compromise interpretability and limit our ability to enforce domain-specific constraints. In the next section, we comprehensively discuss our results with PIP-Loco.
@@ III Results
This section details the simulation and hardware experiments with PIP-Loco. We also present ablation studies on the internal model to validate the importance of a $H$ step Dreamer. For training, we utilized an open-source environment for the Unitree Go1 quadruped robot, which is based on Nvidia’s Isaac Gym simulator [29] [13], and implemented all neural networks using the PyTorch framework [30]. The training process was executed on a desktop system with an AMD Ryzen Threadripper Pro 5995WX processor (128 cores) running at 2.7 GHz, 64 GB of RAM, and an NVIDIA RTX 3080 GPU. For real-time planning, we deployed an onboard computer using Jax [31], achieving an inference frequency of 300-500 Hz.
TABLE III: Performance comparison across different methods and noise levels after 2k iterations for five seeds
@@ III-A Abalations and Experiments
We conducted parallel training of 4096 robots across various terrains, employing domain randomization and a curriculum learning strategy. The PPO algorithm [20] used a clipping range of 0.2, a generalized advantage estimation factor of 0.95, and a discount factor of 0.99. All networks were optimized using the Adam optimizer [32] with a learning rate 0.001.
Figure 5: Simulation results for a 10-second run with extreme command velocities: Top-left compares roll and pitch with and without planning. The top-right shows optimized velocity commands from Algorithm 1. The bottom panel shows joint angle variations, with planning having lower peaks near constraint bounds.
@@ III-A2 What Constitutes a Good Internal Model?
As discussed in [9], an internal model was introduced to simulate a single-step disturbance response in latent space. This work extends that approach by incorporating dreaming of future observations over $H$ steps. Figure 3 illustrates the reward curves for both the model proposed in [9] and our proposed internal models. Additionally, we investigate the influence of the prediction horizon $H$, the model’s robustness to noise, and the role of latent states within our framework. The corresponding results are summarized in Tables III(a) and III(b). Our experimental results suggest that the inclusion of latent states does not improve the robustness of the expert actor $\pi_{e}$ in the presence of training noise. We also observed that base velocity estimation is critical to the overall performance for providing accurate dynamics feedback.
@@ III-A3 Advantages of Planning
We conducted an experiment to assess the robot’s ability to track extreme velocity commands, which can result in unstable behaviors. Fig. 5 presents a comparative analysis between the cloned policy $\pi_{\theta}$, which operates without planning, and Algorithm 1, which incorporates planning. The experiment aimed to track a target velocity of $\text{v}_{x}^{cmd}=2.0$\ \mathrm{m}\mathrm{/}\mathrm{s}$,\text{v}_{y}^{cmd}=2.0$\ \mathrm{m}\mathrm{/}\mathrm{s}$,\omega_{z}^{cmd}=0.5$\ \mathrm{rad}\mathrm{/}\mathrm{s}$$. The planning mechanism in PIP-Loco successfully optimized actions and velocity commands to minimize orientation deviations and maintain base stability while ensuring that joint positions have lower peaks that hit the constraint limits. As a future work, we aim to integrate obstacle avoidance and navigation-related constraints in the planner. The experimental results demonstrate the need for planning in online adaptability. It is important to note that by integrating planning, we could add/remove constraints and reformulate reward/cost functions in real time without retraining. This flexibility is a key characteristic of MPC approaches, which has been notably absent in RL methods for legged locomotion.
@@ IV Conclusion
We introduce PIP-Loco, a novel proprioceptive infinite-horizon planning framework for quadrupedal robot locomotion. Our approach overcomes the limitations of RL-based controllers in terms of interpretability and safety while addressing the task complexity and finite-horizon issue typical of classical MPC methods. By integrating our proposed internal model, PIP-Loco enhances the exploration of effective locomotion behaviors during training and solves an MPC problem at deployment. Extensive ablation studies and multi-terrain tests in both simulation and hardware confirm its robustness and real-world potential.
However, a key limitation is the reliance on the assumption that the expert RL policy and the sampling-based planner behave similarly. Significant deviations during deployment may cause the internal model trained on expert actor behavior to generate unreliable predictions. Future work will explore the integration of learning-based safety filters to align the planner with policy behavior better, as well as incorporate advancements in perceptive locomotion [15] and mitigate dynamics prediction errors.